# Introspective Growth: Automatically Advancing LLM Expertise in Technology Judgment

### v2

@@ Abstract
Large language models (LLMs) increasingly demonstrate signs of conceptual understanding, yet much of their internal knowledge remains latent, loosely structured, and difficult to access or evaluate. We propose self-questioning as a lightweight and scalable strategy to improve LLMs’ understanding, particularly in domains where success depends on fine-grained semantic distinctions. To evaluate this approach, we introduce a challenging new benchmark of 1.3 million post-2015 computer science patent pairs, characterized by dense technical jargon and strategically complex writing. The benchmark centers on a pairwise differentiation task: can a model distinguish between closely related but substantively different inventions? We show that compared to placebo scientific information, prompting LLMs to generate and answer their own questions - targeting the background knowledge required for the task - significantly improves performance. These self-generated questions and answers activate otherwise underutilized internal knowledge. Allowing LLMs to retrieve answers from external scientific texts further enhances performance, suggesting that model knowledge is compressed and lacks the full richness of the training data. We also find that chain-of-thought prompting and self-questioning converge, though self-questioning remains more effective for improving understanding of technical concepts. Notably, we uncover an asymmetry in prompting: smaller models often generate more fundamental, more open-ended, better-aligned questions for mid-sized models than large models with better understanding do, revealing a new strategy for cross-model collaboration. Altogether, our findings establish self-questioning as both a practical mechanism for automatically improving LLM comprehension, especially in domains with sparse and underrepresented knowledge, and a diagnostic probe of how internal and external knowledge are organized. We open-source our patent dataset at this URL.
@@ 1 Introduction
Large language models (LLMs) are increasingly understood not merely as passive text generators, but as systems capable of reasoning about and engaging with complex realities. Recent studies have highlighted their potential for self-improvement—forming hypotheses, seeking clarification, and revising beliefs through iterative interactions [25, 46, 51, 61, 60]. Many of the tasks LLMs perform today demand more than simple fact retrieval. Yet, despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions [4, 19, 41]. Much of their internal knowledge remains latent, poorly organized, and difficult to access [40, 24]. This reflects a fundamental tension: while LLMs appear to “know everything,” their knowledge is stored in a heavily compressed form that impairs activation, recall, and information quality assessment [23, 24, 26]. In this paper, we ask whether LLMs can enhance their own understanding by acting more like learners: generating and answering their own questions – often tailored to their specific uncertainties and usage, reflecting personalized internal gaps rather than universally missing facts – to identify and fill in the background knowledge required for a task. This simple strategy of self-questioning, especially when paired with external retrieval, can automatically improve judgment, expose internal knowledge gaps, and activate reasoning processes that not only enhance the model’s own performance but also transfer effectively to larger models, where such reasoning often remains dormant under standard prompting.
@@
Our first contribution is the proposal of a series of behavioral experiments to systematically evaluate self-questioning as a core mechanism for active learning and introspective growth in LLMs. We focus on the model’s understanding of technology concepts, operationalized through a challenging pairwise differentiation task. Instead of assessing understanding through conventional summarization or question-answer (QA) tasks, we present LLMs with pairs of semantically similar but conceptually distinct patent abstracts and ask whether they describe the same invention. A model that truly understands the material must go beyond surface-level similarity and detect subtle functional or mechanistic differences. To rigorously test this ability, we construct a new dataset of over 1.3 million close-but-different computer science patent pairs, granted post-2015 by the United States Patent and Trademark Office (USPTO), which is our second contribution. Patents offer a uniquely demanding testbed: they are written in dense legal-technical language, often strategically obfuscated, and their distinction is adversarially verified by expert patent examiners. Our task requires fine-grained conceptual discrimination that goes well beyond pattern matching, skills, imitation, or memorization that could possibly bypass the existing benchmarks [36, 64, 12]  [^1] We have more detailed discussions in Appendix A and B.. We embed a structured self-questioning phase between initial exposure and final prediction. Given a pair of closely related patents, the model generates questions targeting both surface-level details and deeper conceptual understanding. It retrieves relevant scientific content to answer its own questions, combining external evidence with internal knowledge. These QA pairs, without verified correctness, are then fed back into the model to inform its final judgment on whether the patents describe the same invention. This setup allows us to assess whether and how self-generated inquiry and targeted retrieval enhance the model’s conceptual understanding (i.e., its accuracy).
@@
Using our framework, we show that prompting LLMs with self-questioning significantly improves their performance on the patent differentiation task, enhancing judgment accuracy, confidence, and response consistency. We then explore three hypotheses to understand what drives this improvement. One possibility is that the questions themselves help the model parse the patent text more effectively by serving as informative cues. Another is that LLMs possess relevant internal knowledge but often fail to access it. A third hypothesis is that even when the model does retrieve relevant knowledge, its internal representations may compress and lose useful information, such that access to external science demonstrably improves performance. To test these possibilities, we design experiments to explore how LLMs’ internal and external knowledge are organized and transferred across model scales, especially when they are sparsely distributed like science, which is our third contribution. We compare the model’s accuracy when given only the questions, when using its questions and self-generated answers, and when using questions and answers derived from science. We find that prompting the model to ask itself questions helps it parse the patent text and improves performance. Feeding LLMs with their self-produced answers also helps, suggesting that LLMs often fail to retrieve relevant internal knowledge. We also find that self-generated answers are less effective than externally sourced ones, however, indicating that hidden compressed capacity remains insufficient for complex judgment tasks, where raw, external scientific knowledge in the form of QA proves beneficial to add richness. We draw a comparison to chain-of-thought (CoT) prompting and find that while they converge, CoT alone fails to recover critical details. This suggests that LLMs lack robust strategies for probing their own internal knowledge. Finally, we investigate how model size influences this process. Interestingly, we find that small models often generate more useful, open-ended, and better-aligned questions for mid-sized models than larger models do, which highlights the importance of matching question complexity to model capacity, as well as a scalable prompting strategy where smaller models help structure the problem in a straightforward way for stronger models to reason through.
@@ 2 A Unique Dataset of Hard-to-Distinguish Patents
Understanding is a complex cognitive ability that goes beyond recalling facts or executing procedures—it involves grasping the deep relational structures that connect concepts, including causal links, analogies, and hierarchies. Drawing from cognitive science and philosophy, a central aspect of understanding is differentiation: recognizing what makes two concepts distinct. This ability is crucial for constructing models of conceptual relationships and is especially vital for causal reasoning, where detecting contrasts and similarities supports counterfactual thinking and explanatory modeling. [^2] Section A in the appendix offers a more detailed discussion of the link between understanding and pairwise differentiation.
Because direct questions like “Do you understand X?” often elicit superficial or memorized responses, we need a more robust test of understanding. Instead, we propose a task that evaluates the model’s capacity to discriminate between concepts that are linguistically similar but semantically distinct. Our test centers on pairwise differentiation: the model is presented with two short texts and asked whether they describe the same underlying idea or two distinct ones. These texts are carefully selected to be semantically similar in language and structure — making the task deceptively difficult, and surface-level heuristics ineffective. The model must detect subtle differences in purpose, mechanism, and/or function, and reason about their implications.
We apply this task in the technology domain and focus specifically on patent abstracts. We construct a novel dataset consisting of 1,319,184 pairs of information and computer technology patent abstracts granted by the USPTO after 2015. In the Appendix C, we also explore another type of "merging" task: whether a model can recognize a rewritten description referring to the same invention. We find that the differentiation task requires genuine conceptual linking and functional comparison — capabilities that cannot be achieved simply by decomposing the task into smaller steps (e.g., via CoT reasoning) or matching texts heuristically – making it a strong testbed for evaluating LLM understanding in the technology domain. Several features make technology patents particularly well suited for testing fine-grained semantic distinctions:
@@
First, patents often describe components of larger systems and therefore exist in dense clusters of near-duplicate-documents that are nearly identical in structure and language but differ in one or two meaningful ways, such as a specific constraint, use case, or technical improvement [53]. These subtle distinctions have been rigorously evaluated by human judges (we will develop this in the third point) and represent precisely the type of contrasts that test whether an LLM can move beyond surface-level similarity and recognize functional difference. We illustrate this pattern in Figure 1(b), which shows the distribution of the most similar—but non-identical—patent pairs based on textual embeddings from [15], pretrained on patent-citation relations. The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart.
Second, patent documents are often strategically obfuscated, written to disclose the minimum required for legal protection while revealing as little as possible about the underlying invention. As a result, their language tends to be intentionally dense, formal, and legally contorted. Readability studies show that patent specifications—particularly the claims sections—require significantly higher grade-level comprehension than comparably technical scientific prose, due to their use of dense jargon and legalistic drafting norms [29, 39]. This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper abstracts from ArXiv and 4,000 matched patent abstracts from the same years. Perplexity, while not a direct measure of understanding, reflects the model’s inability to predict the next word in a sequence. Consistently higher perplexity on patent abstracts suggests that they contain highly technical and domain-specific language that poses a greater challenge for LLMs to parse and comprehend. An example of paired patent abstracts and a related paper abstract can be seen in Appendix D.
Third, and most importantly, even though patents may seem similar, we know that they are not. Each granted patent in our dataset has undergone an extensive expert evaluation process conducted by patent examiners at the USPTO. This review assesses whether the application satisfies the legal requirements of novelty, non-obviousness, and utility. During this process, examiners conduct detailed prior art searches, compare the proposed invention to existing patents and publications, and determine whether the claimed invention constitutes a meaningful and non-trivial contribution over what already exists. This institutional validation is crucial for our task: it provides a strong, domain-informed guarantee that two different granted patents indeed describe conceptually distinct inventions.
@@ 3 Evaluating LLMs’ Understanding of the Technology Domain
We now describe how we operationalize the pairwise differentiation task. We begin by sampling pairs of patents that are difficult to distinguish, identified by embedding all patent abstracts using PATCER [15], a contrastive learning model trained on patent citation relationships. For each patent, we select its most similar—but non-identical—counterpart based on cosine similarity in the embedding space. These pairs are then presented to the LLM, which is asked whether the two abstracts describe the same invention. [^3] See Appendix E for full prompts. For each prompt, the model returns two outputs: a binary label indicating “yes” (1) or “no” (0), and a confidence score on a scale from 0 to 10.
To construct our main measure of understanding, we aggregate predictions using a confidence-weighted voting scheme applied over three independent generations of the same prompt. If the total confidence associated with label “1” exceeds that of label “0,” the final label is set to 1; otherwise, it is 0. This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting [14]. In addition to providing more reliable final decisions, confidence scores offer a fine-grained signal that helps distinguish between borderline and clear-cut cases. This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain [14]. Alongside final labels, we also track changes in confidence and measure self-consistency, defined as the proportion of times all three responses yield the same answer. Lower consistency indicates greater uncertainty in model reasoning.
@@
In the appendix G.1, we show that across all LLaMa models, responses shift systematically depending on how the question is framed. All models are more likely to judge the patents as different when asked explicitly whether they are different, even though the content of the question is logically equivalent. We also report additional results for other model families with varying sizes, including Qwen, Mixtral, and GPT-4o-mini. Again, we find that larger models perform better on the differentiation task. Interestingly, we observe consistent response biases in some models: Mixtral models tend to answer “yes” regardless of the question’s phrasing, while Qwen models tend to answer “no.” Based on these findings, we select the LLaMA-3.1-8B Instruct model as the primary focus for our subsequent experiments and use the question “Do you think they are the same patent?” as the default formulation for the pairwise differentiation task.
@@ 4 Self-questioning LLMs for Improved Technology Understanding
We begin by examining whether prompting models to ask themselves structured questions can enhance performance and, if so, how. We design two types of self-generated questions. The first, surface-level questions, target conceptual recall (e.g., What is X? What does Y do?), while the second encourages deeper understanding through questions about the integration of concepts, improvement, and the intent underlying invention. We construct the pipeline illustrated in Figure 3. When the model is exposed to a pair of closely related patents, $P_{1}$ and $P_{2}$, it generates six questions $Q_{ij}$ for each, where $i\in{1,2}$ indexes the patent and $j\in{1,\dots,6}$ indexes the question. Three of these are surface-level questions, and three target deeper comprehension. These questions guide the retrieval of relevant scientific knowledge, mimicking how researchers ask clarifying questions to deepen understanding. For external knowledge, we use parsed, pre-processed full texts from arXiv computer science papers. Using Google Scholar, we identify the top 10 papers relevant to the patents, segment each into overlapping 2500-character chunks (with 200-character overlaps), and treat these as candidate knowledge units. We embed both the questions and text chunks using the SPECTER2 model [50] specifically for scientific documents pretrained on citation relations, compute cosine similarity, and select the top three relevant chunks per question. These are then passed to the LLM as external knowledge $Knowl_{ij}$. The model answers its own questions using both its internal knowledge and the retrieved content, producing $A_{ij}=f_{LLM}(Q_{ij}\mid Knowl_{ij})$. Then the LLM makes a decision by $Decision=Judge(Patent_{1},Patent_{2}\mid\{(Q_{1j},A_{1j})\}_{j=1}^{6},\{(Q_{2j},A_{2j})\}_{j=1}^{6})$ with the previously mentioned confidence-weighted voting schema.

### v3

# Automatically Advancing LLM Expertise in Technology Judgment

@@ Abstract
Large language models (LLMs) are rapidly becoming core tools for science, engineering, and innovation. Their promise lies not just in remembering facts, but in putting knowledge to work. Despite their impressive ability to answer increasingly difficult questions, it remains unclear whether LLMs truly use their knowledge when confronted with new and challenging tasks. We address this question with a patent classification task that requires deep conceptual understanding: distinguishing objectively different but semantically similar patents. To evaluate this approach, we introduce a challenging new benchmark of 1.3 million post-2015 computer science patent pairs, characterized by dense technical jargon and strategically complex writing. We find that LLMs often fail our benchmark and struggle to distinguish among semantically similar patents. To probe this failure, we introduce a novel framework that decomposes model errors into two sources: missing and unused knowledge. Our approach asks models to generate clarifying questions to improve their understanding, and then compares three settings: raw performance, self-answered questions, and externally supplied answers. This decomposition reveals that LLMs often possess the relevant knowledge internally but fail to deploy it, while a smaller share of errors arises from genuine knowledge gaps. We then ask whether the ability of models to construct a task-specific database of questions and answers differs across models. We find that smaller models generate simpler, broadly transferable questions, while larger models propose more complex but less generalizable ones. This suggests new strategies for combining strengths across models. Taken together, our findings highlight a critical limitation of current LLMs and their evaluation: models often know more than they can use. By shifting evaluation from recall of static facts to application of dynamic knowledge, our approach provides a more informative lens on model capabilities and opens a path toward building systems that better support technological discovery and innovation.
@@ 1 Introduction
Large language models (LLMs) are increasingly understood not merely as passive text generators, but systems capable of reasoning about and engaging with complex realities. Recent studies have highlighted their potential for self-improvement—forming hypotheses, seeking clarification, and revising beliefs through iterative interactions (Huang et al., 2023; Qu et al., 2024; Song et al., 2024; Wang et al., 2023c, b). Yet, despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023). Much of their internal knowledge remains latent, poorly organized, and difficult to access (Pan et al., 2025; Hoang et al., 2023). This reflects a fundamental tension: while LLMs appear to “know everything,” their knowledge is stored in a heavily compressed form that impairs activation, recall, and information quality assessment (Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024).
In this paper, we ask whether LLMs can enhance their own understanding by acting more like learners: generating and answering their own questions – often tailored to their specific uncertainties and usage, reflecting personalized internal gaps rather than universally missing facts – to identify and fill in the background knowledge required for task completion. This simple strategy of self-questioning, especially when paired with external retrieval, can automatically improve judgment, expose internal knowledge gaps, and activate reasoning processes that not only enhance the model’s own performance but also transfer effectively to larger models, where such reasoning often remains dormant under standard prompting.
A further challenge we highlight is the distinction between lay-in knowledge and working knowledge. In humans, once a fact is known, it is typically accessible for reasoning and problem solving: retrieval and use are closely aligned. In LLMs, however, the presence of knowledge in parameters does not guarantee its availability in practice. Prior work has probed this issue by designing tasks that isolate narrow domains (e.g., math word problems or factual recall) to demonstrate what knowledge LLMs are or are not using (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023). By contrast, our framework is situated in a broad technological domain, where the challenge is not the absence of information but the difficulty of activating and applying it in context. This generalizable setup, built on a large and diverse patent corpus, goes beyond special-case demonstrations. It allows us to investigate systematically how LLMs bridge the gap between latent stored knowledge and the working knowledge required for fine-grained conceptual judgment.
@@
Our first contribution is the proposal of a series of behavioral experiments to systematically evaluate self-questioning as a core mechanism for active learning and introspective growth in LLMs. We focus on the model’s understanding of technology concepts, operationalized through a challenging pairwise differentiation task. Instead of assessing understanding through conventional summarization or question-answer (QA) tasks, we present LLMs with pairs of semantically similar but conceptually distinct patent abstracts and ask whether they describe the same invention. A model that truly understands the material must go beyond surface-level similarity and detect subtle functional or mechanistic differences. To rigorously test this ability, we construct a new dataset of over 1.3 million close-but-distinct computer science patent pairs, granted post-2015 by the United States Patent and Trademark Office (USPTO), which is our second contribution. Patents present a uniquely demanding testbed: they are written in dense, legal-technical language, often strategically obfuscated, and their distinctions are adversarially verified by expert patent examiners. Our task requires fine-grained conceptual discrimination that goes well beyond pattern matching, skills, imitation, or memorization that could possibly bypass the existing benchmarks (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023)  [^1] We have more detailed discussions in Appendices A and B.. We embed a structured self-questioning phase between initial exposure and final prediction. Given a pair of closely related patents, the model generates questions targeting both surface-level details and deeper conceptual understanding. It retrieves relevant scientific content to answer its own questions, combining external evidence with internal knowledge. These QA pairs, without verified correctness, are then fed back into the model to inform its final judgment on whether the patents describe the same invention. This setup allows us to assess whether and how self-generated inquiry and targeted retrieval enhance the model’s conceptual understanding (i.e., its accuracy).
@@
Using our framework, we show that prompting LLMs with self-questioning significantly improves their performance on the patent differentiation task, enhancing judgment accuracy, confidence, and response consistency. We then explore three hypotheses to understand what drives this improvement. One possibility is that the questions themselves help the model parse the patent text more effectively by serving as informative cues. Another is that LLMs possess relevant internal knowledge but often fail to access it. A third hypothesis is that even when the model does retrieve relevant knowledge, its internal representations may compress and lose useful information, such that access to external science demonstrably improves performance. To test these possibilities, we design experiments to explore how LLMs’ internal and external knowledge are organized and transferred across model scales, especially when they are sparsely distributed like science, which is our third contribution. We compare the model’s accuracy when given only the questions, when using its questions and self-generated answers, and when using questions and answers derived from science. We find that prompting the model to ask itself questions helps it parse the patent text and improves performance. Feeding LLMs with their self-produced answers also helps, suggesting that LLMs often fail to retrieve relevant internal knowledge. We also find that self-generated answers are less effective than externally sourced ones, however, indicating that hidden compressed capacity remains insufficient for complex judgment tasks, where raw, external scientific knowledge in the form of QA proves beneficial to add richness. Interestingly, we find that small models often generate more useful, open-ended, and better-aligned questions for mid-sized models than larger models do, which highlights the importance of matching question complexity to model capacity, as well as a scalable prompting strategy where smaller models help structure the problem in a straightforward way for stronger models to reason through.
@@ 2 A Unique Dataset of Hard-to-Distinguish Patents
Understanding is a complex cognitive ability that goes beyond recalling facts or executing procedures. It involves grasping the deep relational structures that connect concepts, including causal links, analogies, and hierarchies. Drawing from cognitive science and philosophy, a central aspect of understanding is differentiation: recognizing what makes two concepts distinct. This ability is crucial for constructing models of conceptual relationships and is especially vital for causal reasoning, where detecting contrasts and similarities supports counterfactual thinking and explanatory modeling. [^2] Section A in the appendix offers a more detailed discussion of the link between understanding and pairwise differentiation.
Because direct questions like “what is X?” often elicit superficial or memorized responses, we need a more robust test of understanding. Instead, we propose a task that evaluates the model’s capacity to discriminate between concepts that are linguistically similar but semantically distinct. Our test centers on pairwise differentiation: the model is presented with two short texts and asked whether they describe the same underlying idea or two distinct ones. These texts are carefully selected to be semantically similar in language and structure — making the task deceptively difficult, and surface-level heuristics ineffective. The model must detect subtle differences in purpose, mechanism, and/or function, and reason about their implications.
We apply this task in the technology domain and focus specifically on patent abstracts. We construct a novel dataset consisting of 1,319,184 pairs of information and computer technology patent abstracts granted by the USPTO after 2015. In the Appendix C, we also explore another type of "merging" task: whether a model can recognize a rewritten description referring to the same invention. We find that the differentiation task requires genuine conceptual linking and functional comparison, making it a strong testbed for evaluating LLM understanding in the technology domain. Several features make technology patents particularly well-suited for testing fine-grained semantic distinctions:
@@
First, patents often describe components of larger systems and therefore exist in dense clusters of near-duplicate-documents that are nearly identical in structure and language but differ in one or two meaningful ways, such as a specific constraint, use case, or technical improvement (Trippe, 2015). These subtle distinctions have been rigorously evaluated by human judges (we will develop this in the third point) and represent precisely the type of contrasts that test whether an LLM can move beyond surface-level similarity and recognize functional difference. We illustrate this pattern in Figure 1(b), which shows the distribution of the most similar—but non-identical—patent pairs based on textual embeddings from (Ghosh et al., 2024), pretrained on patent-citation relations. The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart.
Second, patent documents are often strategically obfuscated, written to disclose the minimum required for legal protection while revealing as little as possible about the underlying invention. As a result, their language tends to be intentionally dense, formal, and legally contorted. Readability studies show that patent specifications—particularly the claims sections—require significantly higher grade-level comprehension than comparably technical scientific prose, due to their use of dense jargon and legalistic drafting norms (Kong et al., 2023; Ouellette, 2012). This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper abstracts from ArXiv and 4,000 matched patent abstracts from the same years. Perplexity, while not a direct measure of understanding, reflects the model’s inability to predict the next word in a sequence. Consistently higher perplexity on patent abstracts suggests that they contain highly technical and domain-specific language that poses a greater challenge for LLMs to parse and comprehend. An example of paired patent abstracts and a related paper abstract can be seen in Appendix D.
Third, and most importantly, even though patents may seem similar, we know they are not. Each granted patent in our dataset has undergone an extensive expert evaluation process conducted by patent examiners at the USPTO. This review assesses whether the application satisfies the legal requirements of novelty, non-obviousness, and utility. During this process, examiners conduct detailed prior art searches, compare the proposed invention to existing patents and publications, and determine whether the claimed invention constitutes a meaningful and non-trivial contribution over what already exists. This institutional validation is crucial for our task: it provides a strong, domain-informed guarantee that two different granted patents indeed describe conceptually distinct inventions.
@@ 3 Evaluating LLMs’ Understanding of the Technology Domain
We now describe how we operationalize the pairwise differentiation task. We begin by sampling pairs of patents that are difficult to distinguish, identified by embedding all patent abstracts using PATCER (Ghosh et al., 2024), a contrastive learning model trained on patent citation relationships. For each patent, we select its most similar—but non-identical—counterpart based on cosine similarity in the embedding space. These pairs are then presented to the LLM, which is asked whether the two abstracts describe the same invention. [^3] See Appendix E for full prompts. For each prompt, the model returns two outputs: a binary label indicating “yes” (1) or “no” (0), and a confidence score on a scale from 0 to 10.
To construct our main measure of understanding, we aggregate predictions using a confidence-weighted voting scheme applied over three independent generations of the same prompt. If the total confidence associated with label “1” exceeds that of label “0,” the final label is set to 1; otherwise, it is 0. This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting (Geng et al., 2024). In addition to providing more reliable final decisions, confidence scores offer a fine-grained signal that helps distinguish between borderline and clear-cut cases. This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain (Geng et al., 2024). Alongside final labels, we also track changes in confidence and measure self-consistency, defined as the proportion of times all three responses yield the same answer. Lower consistency indicates greater uncertainty in model judgment.
@@
In the appendix F.1, we show that across all LLaMa models, responses shift systematically depending on how the question is framed. All models are more likely to judge the patents as different when asked explicitly whether they are different, even though the content of the question is logically equivalent. We also report additional results for other model families with varying sizes, including Qwen, Mixtral, and GPT-4o-mini. Again, we find that larger models perform better on the differentiation task. Interestingly, we observe consistent response biases in some models: Mixtral models tend to answer “yes” regardless of the question’s phrasing, while Qwen models tend to answer “no.” Based on these findings, we select the LLaMA-3.1-8B Instruct model as the primary focus for our subsequent experiments and use the question “Do you think they are the same patent?” as the default formulation for the pairwise differentiation task.
@@ 4 Self-questioning LLMs for Improved Technology Understanding
We begin by examining whether prompting models to ask themselves structured questions can enhance performance and, if so, how. We design two types of self-generated questions. The first, surface-level questions, target conceptual recall (e.g., What is X? What does Y do?), while the second encourages deeper understanding through questions about the integration of concepts, improvement, and the intent underlying invention. We construct the pipeline illustrated in Figure 3. When the model is exposed to a pair of closely related patents, $P_{1}$ and $P_{2}$, it generates six questions $Q_{ij}$ for each, where $i\in{1,2}$ indexes the patent and $j\in{1,\dots,6}$ indexes the question. Three of these are surface-level questions, and three target deeper comprehension. These questions guide the retrieval of relevant scientific knowledge, mimicking how researchers ask clarifying questions to deepen understanding. For external knowledge, we use parsed, pre-processed full texts from arXiv computer science papers. Using Google Scholar, we identify the top 10 papers relevant to the patents, segment each into overlapping 2500-character chunks (with 200-character overlaps), and treat these as candidate knowledge units. We embed both the questions and text chunks using the SPECTER2 model (Singh et al., 2023) specifically for scientific documents pretrained on citation relations, compute cosine similarity, and select the top three relevant chunks per question. These are then passed to the LLM as external knowledge $Knowl_{ij}$. The model answers its own questions using both its internal knowledge and the retrieved content, producing $A_{ij}=f_{LLM}(Q_{ij}\mid Knowl_{ij})$. Then the LLM makes a decision by $Decision=Judge(Patent_{1},Patent_{2}\mid\{(Q_{1j},A_{1j})\}_{j=1}^{6},\{(Q_{2j},A_{2j})\}_{j=1}^{6})$ with the previously mentioned confidence-weighted voting schema.